OpenStudy (anonymous):

If (x + yi) + (2(x + yi) + 3i) = 9, what is x + yi?

OpenStudy (freckles):

combine all your real parts on the left hand side and combine all your imaginary parts on the left hand side

OpenStudy (freckles):

can you try to do that @Mr.Man420 and show me what you get

OpenStudy (freckles):

just working with left hand side only

OpenStudy (freckles):

for now

OpenStudy (freckles):

for example we can write 3x+yi+2(x+3yi)+4i as 3x+yi+2x+6yi+4i as 5x+i(y+6y+4) as 5x+i(7y+4) this is just an example of what I want you to do

OpenStudy (freckles):

And pretend we had my example=10 5x+i(7y+4)=10 and we wanted to find x and y such that this equation is true well 5x=10 and 7y+4=0 we set our reals equal to our reals from both sides and our imaginary equal to our imaginary from both sides in my example 5x=10 gives x=2 and 7y+4=0 gives y=-4/7 so the x+yi satisfying my equation 5x+i(7y+4)=10 is 2-4i/7

OpenStudy (anonymous):

Im not quite sure what you want me to do. you still have imaginary on both sides of your equation?

OpenStudy (freckles):

no 9 is real there is no i next to 9

OpenStudy (anonymous):

ok.

OpenStudy (freckles):

I want you to collect your real parts on the left hand side and then collect your imaginary parts on the left hand side basically I'm trying to get you to write your left hand side in standard form of a complex number that is write it in the form u+vi

OpenStudy (freckles):

If (x + yi) + (2(x + yi) + 3i) = 9, what is x + yi? First step distribute the 2 in front of the (x+yi) x+yi+2x+2yi+3i no you see all the things that have an i in it collect them together like I did in my example and everything that doesn't have an i is real so collect those parts together

OpenStudy (freckles):